# The HLLC-2D method for the computation of two phase flow system ejecta transporting model

### v1

@@ Abstract
This work presents a second-order arbitrary Lagrangian Eulerian (ALE) method for the compressible multiphase particle-flow system with the HLLC-2D Riemann solver. We focus on researching the precise equation to describe the interactions between partible phase and flow phase. The computation of equivalent relation of the momentum and energy between two phases is the key point during the procedure, which can be capable of maintaining the conservation of this system. In the case of particle terms, track their trajectories within the mesh and couple with the element based on the position in the space. Thereafter an ALE method instead of Lagrangian scheme is derived for the discretization of the equation to perform better with the complex motion of particles and flow. We apply the HLLC-2D Riemann solver to substitute the HLLC solver which relaxes the requirement for continuous fluxes along the edge. Meanwhile we propose a method for searching particles and provide a CFL-like condition based on this. Finally, we give some numerical tests to analysis the influence of particles on fluid and get a following effect between two phases. The model and method are validated through numerical tests to show its robustness and accuracy.
Keywords:  Multiphase flow, ejecta model, ALE method, HLLC-2D
@@ 1 Introduction
The compressible multiphase flows in a gas-particle system is one of the most important problems in both nature and engineering systems. Due to the high-complexity and multi-scale natures of the gas-particle system, giving an experimental methods of gas-particle system is an investigation so that the numerical simulation of multiphase flows is necessary in the research of the gas-particle flow.
@@
Generally in the early time, systematic research on two-phase flow began in the 1940s, but it was primarily focused on experimental observations and phenomenological descriptions. The term ’Two-phase Flow’ first appeared in the literature in [15]. After that, people started to focus on this problem which led to the gradual emergence of numerous research papers and monographs on multiphase flow theory. Over the past few years, the study of two-phase flow has evolved from relatively simple pipe flows to more complex fluidized beds and gas-particle two-phase separation flows. It has also progressed from low concentration (dilute particle phase) flows to high concentration (dense particle phase) flows. Building upon the foundation of dilute phase flow, researchers have developed physical models and fundamental equations for dense phase flow[22, 23, 24, 25, 26]. Subsequently, many researchers, including Campbell et al.[27], conducted a series of studies and established the theory of granular dynamics. Then Ding and Gidaspow[28] further developed and refined the theory of granular dynamics, introducing the concept of virtual temperature to characterize the effects of particle collisions in dense granular flows. They also performed numerical simulations that yielded results consistent with experimental observations.
@@
In the recent years, the complex interactions in multiphase flow between phases have garnered increasing attention. Many scholars have conducted detailed research on turbulent disturbances in two-phase flows, with particular emphasis on simulating gas-solid two-phase turbulence, which has become a research focus. Various approaches have been developed, including tow-phase turbulence direct numerical simulation[29], Reynolds-averaged numerical simulation and large-eddy simulation[51, 50].
@@
Especially in two-phase problems, subject to extreme shock loading, metal materials may undergo fragmentation, resulting in high-speed disintegration of particulate matter separating from the metal materials. This particulate matter then mixes with the surrounding gas, leading to a phenomenon called ’ejecat mixing’[34, 35]. The theoretical and experimental method and numerical simulation of ejecta mixing is currently a challenging and prominent issue in the field of implosion compression science [29, 30, 31, 33]. It represents one of the significant challenges in two-phase flow and is also a highly important area of application.
@@
There are many experimental results publishing on the ejecta mixing. In 1983, Elias et al. published the experimental results of jet mixing involving metallic tin and argon gas [36]. They conducted joint observations of ejecta mixing evolution using X-ray imaging and high-speed photography. Later Ogorod.nikov et al. designed ejecta mixing experiments involving metallic lead and air[37]. In the same year, Buttler et al. [39] utilized PDV technology to investigate the velocity evolution of tin ejecta in neon gas. Their experimental results reflected the gradual decay of the ejecta over time due to fragmentation and resistance effects. In [38], Oro et al. conducted ejecta mixing experiments by preloading tungsten particles with a certain size distribution on an aluminum substrate using explosive loading, including vacuum ejecta mixing experiments, argon gas filling experiments, and hydrogen gas filling experiments. Sorenson et al. [40] further studied the characteristics of ejecta size distribution with different gases.
@@
On the other hand, studying particle motion through numerical simulation is also an important role to supplement experiments.In 2004, Wang et al.[52] used SPH method to simulate the ejecta phenomenon on the metal surface under shock loading. Harrison, Fung, and others[11, 41], based on the FLAG Lagrangian fluid dynamics program platform, have developed an initial numerical simulation program for ejecta mixing. This program consists of two main modules: the particle source module and the particle transport module. The particle source module is used to generate and initialize particles, including determining the conditions for ejecta initiation, the total mass of ejecta, the mass and velocity distribution of the ejecta, the size distribution of the particles, the thermodynamic state of the particles, initial velocity, initial position, and more. The particle transport module is employed to simulate the movement of particles in the gas environment, with the main algorithm utilizing the MP-PIC method [44, 45, 12]. In 2013, Fung et al. [42] provided a detailed introduction to the method of introducing initial particles and their state settings, but did not present quantitative calculations of ejecta mixing results or comparisons with experimental data.
@@
Due to the complexity of the coupled fluid-particle system, it means that while the fluid influences the particles, the particles also exert an effect on the fluid, necessitating the use of robust numerical methods for fluid model computations. In recent years, a classical numerical method for fluid equation is primarily based on a conserved cell-centered grid discretization. The one-dimensional scheme of this method was initially proposed by Godunov et al. [46, 13], and later it was extended to multi-dimensional scenarios [47, 48]. And later many scholars have proposed approximate Riemann solvers at the nodes, such as Despres and Mazera [49], Maire [3], which have shown significant improvements in correcting results.
@@
Drawing from the advancements outlined in the above-mentioned literature, we introduce a coupled particle-fluid transporting model by configuring the initial states of the particles in a case of low concentration. And we focus on solving particle-gas system transporting problem based on a cell-centered Arbitrary Lagrangian-Eulerian (ALE) method which is the generalization and application from the ALE frame in [1].
@@
Based on the ejecta mixing and a coupled transporting effect, we apply the particle trajectory model to analyze the interactions and energy exchange between gas phase and particle. Through the numerical tests, we show the motion results of particles and fluid and a following effect of particles in fluid. Furthermore, we analyze the parameters influencing this effect and the reasons for its formation.
@@
The rest of the paper are organized as follows. In Section 2, we present the particle trajectory model and give the governing equations. In Section 3, a numerical simulation method is constructed. In Section 4 we give the time discretization and a particle searching strategy for an extra CFL-like condition to control the time step. In Section 5 we show the numerical tests, analyze the results and give a conclusion.
@@ 2 Particle Trajectory Model
In this section, we shall study the particle trajectory model by analyzing the motion of particle and fluid. It is a two-dimensional model describing the compressible two-phase problem. Though multiphase flow are more complex than single-phase flow, they still follow the basic of mass, momentum and energy conservation in essence. So the model must maintain the conservation of the system.
@@ 2.1 Problem statement, Notation and Assumptions
Next we will introduce an ejecta mixing model considering the following problem. Multiple particles mix in the fluid(the total volume is less than a certain percentage). Regard each particle as an individual, interacting with the fluid separately and explore the motion of each particle independently. Since our particle density is less than a certain range, the particle flow is considered to be sparse particle flow, so it is assumed that each particle does not affect each other. And the fluid has initial density, velocity, pressure and interacts with particles to exchange momentum and energy at the same time.
In the following part, subscript $p$ indicates the physical parameter that is about particle phase, subscript $g$ is about the gas phase. Let $\rho_{g}$, $\bm{v_{g}}=(v_{g_{x}},v_{g_{y}})$, $E_{g}$, $\mu_{g}$, $T_{g}$, $C_{g}$ be the density, velocity, energy, viscosity and thermal conductivity of the fluid and $P$ be the pressure of the space. $\bm{u_{p}}=(x,y)$, $\rho_{p}$, $\bm{v_{p}}=(v_{p_{x}},v_{p_{y}})$, $T_{p}$, $c_{p}$, $r_{p}$ denote the coordinate, density, velocity ,temperature, specific heat capacity and radius of particle. In particular, $\bm{u_{p_{m}}}=(x_{m},y_{m})$, $\rho_{p_{m}}$, $\bm{v_{p_{m}}}$, $T_{p_{m}}$, $c_{p_{m}}$ $r_{p_{m}}$denote the variable of the particle $m$, in which $m=1,\ldots,m(p)$, $m(p)$ is the total number of particles.
To explore the interaction between two phases, the force analysis of the particles in the system is required. During the movement, particles are subject to gravity, drag, differential pressure, Basset force, Magnus force, Saffman force, etc. In general, not all of these forces are equally important. It is necessary to compare the magnitude of the above forces in order to estimate the relative importance of the various forces. Under normal conditions, other forces are several orders of magnitude smaller than the drag. So for the convenience of analysis, above forces except drag are not considered in the following model.
In addition to the force analysis, the transfer of heat between the particles and the fluid is also significant. Normally there are three ways for transferring the heat including heat conduction, convection heat transfer, thermal radiation. Similarly, because of the orders of the magnitude, we only consider the convection heat transfer.
@@
    Consider ideal compressible gas. For each case we use an ideal gas equation of state, namely $P=(\gamma-1)\rho\epsilon$ , where $\gamma$ is the adabatic index.
@@
    Viscosity and heat transfer exist only in interphase interactions.
@@ 2.2 Equations
2.2.1 Governing equation
The governing equations for two-dimensional compressible particle trajectory model are as follows:
$\frac{\partial\rho_{g}}{\partial t}+\nabla\cdot(\rho_{g}\bm{v}_{g})=0,$ (2.1)
$\frac{\partial}{\partial t}(\rho_{g}\bm{v}_{g})+\nabla\cdot(\rho_{g}\bm{v}_{g}\otimes\bm{v}_{g})+\nabla P=-\bm{F}_{p},$
$\frac{\partial}{\partial t}(\rho_{g}E_{g})+\nabla\cdot(\rho_{g}E_{g}\bm{v}_{g})+\nabla\cdot(p\bm{v}_{g})=-{E_{p}},$
$\frac{\partial\bm{u_{p}}}{\partial t}=\bm{v}_{p},$ (2.2)
$\frac{\partial\bm{v}_{p}}{\partial t}=\bm{F}_{p},$
$\frac{4}{3}\pi\,r_{p}^{3}\,\rho_{p}\,c_{p}\frac{\partial T_{p}}{\partial t}=Q_{p},$
where the $\bm{F}_{p}$ and ${E_{p}}$ indicate the exchange of momentum and energy due to interaction between the particles and fluid. In the particle trajectory model, the motion of every particles will be calculated respectively, we have
$\bm{F}_{p}=\sum_{m=1}^{m(p)}\bm{F}_{p_{m}},\$ (2.3)
${E_{p}}=\sum_{m=1}^{m(p)}{E}_{p_{m}}.\$ (2.4)
$\bm{F}_{p_{m}}$ is the force between particle ${}^{\prime}m^{\prime}$ and the fluid which is the source power of the particle’s movement. ${E}_{p_{m}}$ is the exchange of energy including the heat and the kinetic energy between the particle ${}^{\prime}m^{\prime}$ and fluid.
2.2.2 Interaction and Conservation
Next we give the specific definition of interaction items $\bm{F}_{p}$ and ${E_{p}}$.
The force $\bm{F}_{p}$ comes from the drag due to friction. There are two drag models usually used during computing which are the Stoke’s model and Crowe’s model.
@@ 2.2.2 Interaction and Conservation
$\bm{F}_{p}=\frac{1}{2}C_{d}S_{p}\rho_{g}\left|\bm{v}_{g}-\bm{v}_{p}\right|\left(\bm{v}_{g}-\bm{v}_{p}\right),$ (2.6)
where $C_{D}$ is the drag coefficient and $S_{p}$ is the upwind area. Drag coefficient $C_{D}$ is related to the Reynolds number $Re$, the following definition is applied[4][8]:
$C_{D}=\frac{24}{Re_{p}}f_{drag}(Re).$ (2.7)
Function $f_{drag}(Re)$ depends on the Reynolds number $Re$, it has:
$Re=\frac{2\rho_{g}r_{p}\left|\bm{v}_{g}-\bm{v}_{p}\right|}{\mu_{g}},$
@@
Remark: The calculation method of $C_{\mathrm{D}}$ is not same in different research. The above definition is given by Wang et al in paper [8].
The energy interaction consists of two parts of which are drag work and heat transfer, namely:
${E_{p}}=\bm{F}_{p}\cdot\bm{v}_{p}+Q_{p}.$ (2.9)
$Q_{p}$ comes from the convection heat transfer, which is[7]:
$Q_{p}=\frac{2\mu_{g}}{Pr}c_{pg}\pi r_{p}\left(T_{g}-T_{p}\right)Nu_{p},$ (2.10)
where $T_{g}$ and $T_{p}$ is the temperature of fluid and particle respectively. $C_{pg}$ is constant-pressure specific heat capacity, $Nu_{p}$ is Nusselt numbers, $Pr$ is Prandtl numbers of flow, of which specific definitions are as follows:
@@
$Nu_{p}=2+0.6Pr^{1/3}Re^{1/2}.$ (2.12)
Based on the equation (2.1) and (2.2), we need to prove that the above gas-particle two phase system satisfies conservation law. It is obvious that the conservation of momentum because the drag between the particle and the fluid is a pair of interacting forces, namely the right items of the second equation of (2.1) and of the equation (2.2). The item $\bm{F}_{p}\bm{v}_{p}$ is the energy of the fluid which is converted into the kinetic energy of the particle. The convection heat transfer energy $Q_{p}$ is generated by the temperature difference between the fluid and the particle, which is from the higher phase to the lower phase.

### v2

# An ALE numerical method with HLLC-2D solver for the two-phase flow ejecta transporting model

@@ Abstract
This work presents an arbitrary Lagrangian Eulerian (ALE) method for the compressible two-phase flow ejecta transporting model with the HLLC-2D Riemann solver. We focus on researching the precise equation to describe the interactions between particle phase and flow phase. The calculation of the momentum and energy exchange across two phases is the key point during the procedure, which can be capable of maintaining the conservation of this system. For particles, tracking their trajectories within the mesh and elements is essential. Thereafter an ALE method instead of Lagrangian scheme is derived for the discretization of the equation to perform better with the complex motion of particles and flow. We apply the HLLC-2D Riemann solver to substitute the HLLC solver which relaxes the limitation for continuous fluxes along the edge. Meanwhile we propose a method for searching particles and provide a CFL-like condition based on this. Finally, we show some numerical tests to analysis the influence of particles on fluid and get a following effect between two phases. The model and the numerical method are validated through numerical tests to show its robustness and accuracy.
Keywords:  two-phase flow, ejecta transporting model, ALE method, HLLC-2D
@@ 1 Introduction
The compressible multiphase flows in a gas-particle system is one of the most important problems in both nature and engineering systems. Due to the high-complexity and multi-scale natures of the gas-particle system, providing effective experimental methods and numerical simulations is essential in the research of gas-particle flow.
@@
Generally in the early time, systematic research on two-phase flow began in the 1940s, but it was primarily focused on experimental observations and phenomenological descriptions. The term ‘Two-phase Flow’ first appeared in the literature in [40]. After that, people started to focus on this problem which led to the gradual emergence of numerous research papers and monographs on multiphase flow theory. It has also progressed from low concentration (dilute particle phase) flows to high concentration (dense particle phase) flows. Building upon the foundation of dilute phase flow, researchers have developed physical models and fundamental equations for dense phase flow[33, 21, 23, 15, 35]. Subsequently, many researchers, including Campbell et al.[6], conducted a series of studies and established the particle dynamics. Then Ding and Gidaspow[12] further developed and refined the theory of particle dynamics, introducing the concept of virtual temperature to characterize the effects of particle collisions in dense granular flows. They also conducted numerical simulations which yielded results consistent with experimental observations.
@@
In the recent years, the complex interactions in multiphase flow between phases have garnered increasing attention. Many scholars have conducted detailed research on turbulent disturbances in two-phase flows, with particular emphasis on simulating gas-solid two-phase turbulence, which has become a research focus. Various approaches have been developed, including two-phase turbulence direct numerical simulation[32], Reynolds-averaged numerical simulation and large-eddy simulation[7, 4].
@@
Under extreme shock loading, particularly in two-phase problems, metal materials’ surface probably experiences fragmentation, resulting in high-speed particles separating from the metal materials. This particulate matter then mixes with the surrounding gas, leading to a phenomenon called ’ejecat mixing’[43, 2]. The theoretical and experimental method and numerical simulation of ejecta mixing is currently a challenging and important issue in the field of implosion compression science [32, 20, 47, 46]. It represents one of the significant challenges in two-phase flow and is also a highly important area of application.
@@
There are many experimental results publishing on the ejecta mixing. In 1983, Elias et al. published the experimental results of jet mixing involving metallic tin and argon gas [14]. They conducted joint observations of ejecta mixing evolution using X-ray imaging and high-speed photography. Later Ogorodnikov et al. designed ejecta mixing experiments involving metallic lead and air[29]. In the same year, Buttler et al. [5] utilized PDV technology to investigate the velocity evolution of tin ejecta in neon gas. Their experimental results reflected the gradual decay of the ejecta over time due to fragmentation and resistance effects. In [30], Oro et al. conducted ejecta mixing experiments by preloading tungsten particles with a certain size distribution on an aluminum substrate using explosive loading, including vacuum ejecta mixing experiments, argon gas filling experiments, and hydrogen gas filling experiments. Sorenson et al. [39] further studied the characteristics of ejecta size distribution with different gases.
@@
On the other hand, studying particle motion through numerical simulation is also an important role to supplement experiments.In 2004, Wang et al.[44] used SPH method to simulate the ejecta phenomenon on the metal surface under shock loading. Harrison, Fung, and others[16], based on the FLAG Lagrangian fluid dynamics program form, have developed an initial numerical simulation program for ejecta mixing. This program consists of two main models: the particle source model and the particle transport model. The particle source model is used to generate and initialize particles, including determining the conditions for ejecta initiation, namely the mass, the velocity distribution, the size distribution, the thermodynamic state, initial position of the particles and more. The particle transport model is given to simulate the motion of particles in the gas environment, with the main algorithm utilizing the MP-PIC method [1, 37, 36]. In 2013, Fung et al. [16] provided a detailed introduction to the settings of particles’ initial state, but did not present quantitative calculations of ejecta mixing results or comparisons with experimental data.
@@
Because of the complexity of the coupled gas-particle ejecta system, it means that while the fluid influences the particles, the particles also exert an effect on the fluid, necessiating the use of robust numerical methods for fluid motion simulations. In earlier years, a classical numerical method for fluid equation is primarily based on a conserved cell-centered grid discretization. The one-dimensional scheme of this method was initially presented by Godunov et al. [17], and later it was extended to multi-dimensional scenarios [3, 13]. Many scholars have proposed approximate Riemann solvers at the nodes, such as Després and Mazera [11], Maire et al.[24], which have shown significant improvements in correcting results.
@@
Drawing from the advancements outlined in the above-mentioned literature, we employ a coupled gas-particle ejecta transporting model through configuring the initial states of the particles in a case of low concentration. And we focus on solving this problem using a cell-centered Arbitrary Lagrangian-Eulerian (ALE) method which is an extension of the method in [34] from computational single-phase flow to the calculation of two-phase flow problems.
@@
Based on the ejecta mixing and a coupled transporting effect, we apply the particle trajectory model to analyze the interactions and energy exchange between gas and particle phase. Through the numerical tests, we show the motion results of particles and fluid and a following effect of particles in fluid. Furthermore, we analyze the parameters that influence this effect and explore the reasons behind its formation.
@@
The rest of the paper are organized as follows. In Section 2, we introduce the particle trajectory model and interactions between two phases. In Section 3, a numerical simulation method is given. In Section 4, we present the time discretization and a particle searching strategy for a CFL-like condition to limit the time step. In Section 5, we show the numerical tests, analyze the results and give a conclusion.
@@ 2 Particle Trajectory Model
In this section, we introduce the two-dimensional particle trajectory model by analyzing the dynamics of particles and compressible fluid.
@@ 2.1 Problem statement, Notation and Assumptions
Consider the following problem in a region $\Omega$ of two-dimensional space that multiple particles are mixed in the fluid(the total volume is less than a certain percentage). Regard each particle as an individual, interacting with the fluid separately and explore the motion of each particle independently. Since our particle concentration is less than a certain ratio, the particle flow is considered to be sparse particle flow, so it is assumed that each particle does not affect each other. And the fluid has initial density, velocity, pressure and interacts with particles to exchange momentum and energy at the same time.
In the following part, subscript $p$ indicates the physical parameter that is about particle phase, subscript $g$ is about the gas phase. Let $\rho_{g}$, $\bm{v_{g}}=(v_{g_{x}},v_{g_{y}})$, $E_{g}$, $\mu_{g}$, $T_{g}$, $C_{g}$ be the density, velocity, energy, viscosity, thermal conductivity and $P$ be the pressure of the fluid. $\bm{u_{p}}=(u_{p_{x}},u_{p_{y}})$, $\rho_{p}$, $\bm{v_{p}}=(v_{p_{x}},v_{p_{y}})$, $T_{p}$, $c_{p}$, $r_{p}$ denote the coordinate, density, velocity ,temperature, specific heat capacity and radius of particle. In particular, if there are multiple particles in the region $\Omega$ with different physical quantities, we can additionally use a subscript $m$ to identify them. Namely $\bm{u_{p_{m}}}$, $\rho_{p_{m}}$, $\bm{v_{p_{m}}}$, $T_{p_{m}}$, $c_{p_{m}}$ $r_{p_{m}}$denote the variable of the different particles, in which $m=1,\ldots,m(\Omega)$, $m(\Omega)$ is the total number of particles in the region $\Omega$.
To investigate the interactions between two phases, the force analysis of the particles in the system is essential. During the motion, particles are subject to gravity, drag, differential pressure, Basset force, Magnus force, Saffman force, etc. In general, not all of these forces are equally important. It is necessary to compare the magnitude of the above forces in order to estimate the relative importance of the various forces. Under normal conditions, other forces are several orders of magnitude smaller than the drag. So for the convenience of analysis, above forces except drag are not considered in the following model.
In addition to the force analysis, the exchange of heat between the particles and the fluid is also significant. Normally there are three ways for transferring the heat including heat conduction, convection heat transfer, thermal radiation. Similarly, because of the orders of the magnitude, we only consider the convection heat transfer.
@@
    For each case we consider ideal compressible gas with ideal gas equation of state $P=(\gamma-1)\rho\epsilon$, where $\gamma$ is the adabatic index.
@@
    Viscosity and heat transfer exist only in interface interactions.
@@ 2.2 Governing Equations
The equations for two-dimensional compressible particle trajectory model are as follows, for the fluid phase:
where the state vector and flux vectors are
$\mathbf{U}=\left[\begin{array}[]{c}\rho_{g}\\ \rho_{g}v_{g_{x}}\\ \rho_{g}v_{g_{y}}\\ \rho_{g}E_{g}\end{array}\right],\quad\mathbf{F}(\mathbf{U})=\left[\begin{array}[]{c}\rho_{g}v_{g_{x}}\\ P+\rho_{g}v_{g_{x}}^{2}\\ \rho_{g}v_{g_{x}}v_{g_{y}}\\ \rho_{g}E_{g}v_{g_{x}}+Pv_{g_{x}}\end{array}\right],\quad\mathbf{G}(\mathbf{U})=\left[\begin{array}[]{c}\rho_{g}v_{g_{y}}\\ \rho_{g}v_{g_{x}}v_{g_{y}}\\ \rho_{g}v_{g_{y}}^{2}+P\\ \rho_{g}E_{g}v_{g_{y}}+Pv_{g_{y}}\end{array}\right].$ (2.2)
For the particle phase, each particle has:
$\frac{\partial\bm{u_{p_{m}}}}{\partial t}=\bm{v}_{p_{m}},$ (2.3)
$\frac{\partial\bm{v}_{p_{m}}}{\partial t}=\bm{F}_{p_{m}},$
$\frac{4}{3}\pi\,r_{p}^{3}\,\rho_{p}\,c_{p}\frac{\partial T_{p_{m}}}{\partial t}=Q_{p_{m}},$
where the $\bm{F}_{p_{m}}$ and ${Q_{p_{m}}}$ indicate the interactions between the particle phase and fluid phase.
And $\mathbf{S}_{p}$ is the resource term between the flow and particles
$\mathbf{S}_{p}=\left[\begin{array}[]{c}0\\ {F}_{p_{x}}\\ {F}_{p_{y}}\\ {E}_{p}\end{array}\right],$ (2.4)
where $\bm{F}_{p}$ is the force which is the source item of the particle’s motion. ${E}_{p}$ is the exchange of energy including the heat and the kinetic energy.
In the particle trajectory model, the motion of every particles will be calculated respectively, actually we have
$\bm{F}_{p}=\left[\begin{array}[]{c}{F}_{p_{x}}\\ {F}_{p_{y}}\\ \end{array}\right]=\sum_{m=1}^{m(\Omega)}\bm{F}_{p_{m}},\$ (2.5)
${E_{p}}=\sum_{m=1}^{m(\Omega)}{E}_{p_{m}}=\sum_{m=1}^{m(\Omega)}\bm{F}_{p_{m}}\cdot\bm{v}_{p_{m}}+Q_{p_{m}}.\$ (2.6)
Then give the specific definition of interaction items $\bm{F}_{p_{m}}$ and ${Q_{p_{m}}}$. The force $\bm{F}_{p_{m}}$ comes from the drag due to friction. There are two drag models usually used during computing which are the Stoke’s model and Crowe’s model.
@@
$\bm{F}_{p_{m}}=\frac{1}{2}C_{d}K_{p_{m}}\rho_{g}\left|\bm{v}_{g}-\bm{v}_{p_{m}}\right|\left(\bm{v}_{g}-\bm{v}_{p_{m}}\right),$ (2.8)
where $C_{D}$ is the drag coefficient and $K_{p_{m}}$ is the upwind area. Drag coefficient $C_{D}$ is related to the Reynolds number $Re$, the following definition is applied[27][45]:
$C_{D}=\frac{24}{Re}f_{drag}(Re).$ (2.9)
In formulation (2.9), $f_{drag}(Re)$ depends on the Reynolds number $Re$:
$Re=\frac{2\rho_{g}r_{p_{m}}\left|\bm{v}_{g}-\bm{v}_{p_{m}}\right|}{\mu_{g}},$
@@
The calculation method of $C_{\mathrm{D}}$ is not same in different research. The above definition is given by Wang et al in [45].
$Q_{p_{m}}$ comes from the convection heat transfer[8]:
$Q_{p_{m}}=\frac{2\mu_{g}}{Pr}c_{pg}\pi r_{p_{m}}\left(T_{g}-T_{p_{m}}\right)Nu_{p_{m}},$ (2.11)
where $C_{pg}$ is constant-pressure specific heat capacity, $Nu_{p_{m}}$ is Nusselt numbers, $Pr$ is Prandtl numbers of flow, of which specific definitions are as follows:
@@
$Nu_{p_{m}}=2+0.6Pr^{1/3}Re^{1/2}.$ (2.13)
Based on the equation (2.1) and (2.3), we need to prove that the above gas-particle two phase system satisfies conservation law. It is obvious that the conservation of momentum because the drag between the particle and the fluid is a pair of interacting forces, namely the right items of the second equation of (2.1) and of the equation (2.3). The item $\bm{F}_{p_{m}}\bm{v}_{p_{m}}$ is the energy of the fluid which is converted into the kinetic energy of the particle and the convection heat transfer energy $Q_{p_{m}}$ is generated by the temperature difference.